The Townson Manufacturing Company has gathered the following information for the month of September:

  • 11,100 units in the beginning Work-in-Process Inventory (80% complete as to materials, 1/3 complete with respect to the conversion costs).
  • 77,000 units were started into production.
  • 67,000 units were completed and transferred to the next department.
  • The ending Work-in-Process Inventory is complete as to materials but only 3/8 complete with respect to conversion costs.

What are the equivalent units of production (EUPs) for materials in the month of September assuming Townson uses weighted-average process costing?

Multiple Choice

  • 70,700 EUPs

  • 85,880 EUPs

  • 88,100 EUPs

  • 79,220 EUPs
Answer #1

answer:88100EUPs

   Equivalent units of production for material

Beginning work in process+units were started in production-Units completed and transferred to next department= ending inventory

Ending inventory= 11100+77000-67000=21100

   EQUIVALENT UNIT OF PRODUCTION for materials(weighted average method)

Units completed and transferred to next department= 67000×100%completion=67000

Ending inventory= 21100×100%completion= 21100

Total equivalent units of production for materials = 88100EUPs

     

   ALL THE BEST

PLEASE DO SUPPORT US

ANY doubt please comment BELOW

Thank you